# Computer-Assisted Proofs of Solitons in Bose-Einstein Condensates

### v1

@@ Abstract
We rigorously prove the existence of gap solitons in the one-dimensional Gross-Pitaevskii (GP) equation with a periodic potential. These nonlinear localized solutions emerge in spectral gaps and play a crucial role in understanding Bose-Einstein condensates (BECs). To prove them, we reformulate the problem as the search for homoclinic orbits in a higher-dimensional dynamical system. We then use computer-assisted proof techniques, combined with a functional analytic framework, to rigorously validate numerically approximated homoclinic orbits. This work bridges computational evidence and formal mathematical proofs, providing a solid foundation for the study of solitons in the GP equation.
Key words. Bose-Einstein condensates, Gross-Pitaevskii equation, Gap solitons, Homoclinic orbits, Conservative systems, Parameterization method for periodic orbits, Computer-Assisted Proofs
@@ 1 Introduction
A Bose-Einstein condensate (BEC) is a state of matter that forms when a collection of particles cools to temperatures near absolute zero, causing them to lose their individual identities and merge into a single quantum wave. The first experimental realization of BECs at ultra-cold temperatures earned the 2001 Nobel Prize in Physics, marking a major breakthrough in science. Beyond their fundamental significance, BECs provide a powerful platform for exploring quantum mechanics on large scales, with far-reaching applications in precision measurements, quantum computing, and the modeling of complex physical systems such as superfluidity and optical lattices.
The mathematical richness of Bose-Einstein condensates offers fertile ground for advancing theoretical methods in nonlinear dynamics and partial differential equations. The book [1] serves as a comprehensive resource on nonlinear wave dynamics, combining experimental insights with numerical analyses of BECs. It also includes an extensive bibliography with thousands of references spanning various aspects of the field.
In this paper, we study the dynamics of a BEC under the mean-field approximation, where the condensate is described by a macroscopic wave confined in an elongated, cigar-shaped trap and subjected to a one-dimensional optical lattice along the axis of weaker confinement. We express the system’s energy in terms of the condensate’s density and phase variables, leading to the time-dependent Gross-Pitaevskii equation, which governs the BEC’s evolution in one spatial dimension:
@@
Here, $\psi(t,x)\in\mathbb{C}$ denotes the dimensionless wave function, $|\psi|^{2}$ represents the BEC density, and $V(x)$ is the external potential created by the optical lattice along the spatial domain $x\in\mathbb{R}$. The sign of $c$ determines the interaction behavior between particles in the condensate: for $c>0$, interactions exhibit self-defocusing behavior, while for $c<0$, they become self-focusing. Both regimes are physically significant. A comprehensive discussion, covering both experimental and theoretical perspectives, can be found in [2].
A fundamental aspect of BEC analysis is the study of standing wave solutions, which take the form $\psi(t,x)=e^{-iat}u(x),$ where $u(x)$ is a real-valued function that satisfies the time-independent equation
@@
We refer to this equation as the Gross-Pitaevskii (GP) equation, where the parameter $a$ represents the wave frequency.
The GP equation admits a variety of nonlinear structures, such as solitons, nonlinear Bloch waves and domain walls, which have been extensively studied in the literature (e.g. see [2, 3]). In particular, solitons are a class of localized solutions $u:\mathbb{R}\to\mathbb{R}$ satisfying the additional asymptotic condition
@@
Chapter 19 in [1] provides a general discussion and references on the existence of solitons, depending on different properties of the potential, and particularly focuses on solitons for periodic potentials. In this paper, we focus our study on the periodic potential given by $V(x)=b\cos(2x)$, which describes the dynamics of a Bose-Einstein condensate in a periodic optical lattice.
To better understand the solutions of the GP equation (2), we can first examine the linearized problem around the trivial solution, which leads to Mathieu’s differential equation
$Lu\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,\left(\partial_{x}^{2}+a-V(x)\right)u=\left(\partial_{x}^{2}+a-b\cos(2x)\right)u=0.$ (4)
In this framework, Bloch’s theory predicts bands of continuous spectrum separated by spectral gaps. The boundaries of these gaps are determined by Mathieu functions, which are solutions to (4). In the linear regime defined by the solutions of (4), soliton-type solutions are not possible. However, when nonlinearity is introduced, localized modes of the condensate, known as gap solitons, can form within the spectral gaps. These gap solitons are a hallmark of nonlinear wave phenomena and are analogous to gap solitons observed in nonlinear optics. As discussed in [3], gap solitons manifest in a variety of nonlinear optical settings, including periodic structures, waveguide arrays, and optically induced photonic lattices.
The existence of gap solitons in multiple dimensions has been established through abstract theorems presented in [4, 5, 6, 7] and the references therein, relying on a range of variational techniques. Nevertheless, these results provide no explicit characterization of the solution profiles. The GP equation has been studied in the context of bounded, non-periodic potentials $V(x)$, as in [8], where the nonlinearity differs from the cubic term in (2). More precisely, bifurcations of solitons are shown to originate from eigenvalues located to the left of the continuous spectrum. Conversely, works such as [9] address bifurcations from the boundary of the continuous spectrum, exploring the broader question of solution bifurcations for nonlinear operator equations when the linearized operator has no eigenvalues.
Even in one-dimensional settings, proving the existence of gap solitons while providing explicit characterization of the solution profile is highly non-trivial. Consequently, researchers have frequently relied on perturbative asymptotics and numerical simulations to investigate these solutions. For instance, [3] utilizes asymptotic analytical methods to compute gap solitons in all band gaps of a periodic potential, regardless of the sign of the nonlinearity. These solitons bifurcate from distinct band edges depending on whether $c$ is positive or negative. Moreover, [10] introduces a numerical approach for calculating gap solitons in the case of repulsive interactions ($c=1$). This study demonstrates that most solutions exhibit blow-up behavior at specific points on the real axis. Notably, the subset of non-blow-up solutions, which includes all gap solitons, corresponds to a fractal set within the space of initial conditions. This fractal structure enables the identification of gap solitons over large regions in the parameter space $(a,b)\in\mathbb{R}^{2}$.
Although these simulations and asymptotic techniques provide valuable insights, they often fall short of offering rigorous proof of soliton existence. In this paper, we address this limitation by presenting a methodology that rigorously validates numerically approximated solitons in non-perturbative parameter regimes, thereby bridging the gap between computational simulations and formal mathematical proofs.
@@
More precisely, denote $u_{1}\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,u$, $u_{2}\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,u^{\prime}$, $u_{3}\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,V(x)=\cos(2x)$ (the periodic potential) and $u_{4}\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,u_{3}^{\prime}=-2\sin(2x)$. Note that $u_{3}$ solves $u_{3}^{\prime\prime}=-4u_{3}$ with initial conditions $(u_{3}(0),u_{3}^{\prime}(0))=(1,0)$. Assume that $U\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,(u_{1},u_{2},u_{3},u_{4})\in\mathbb{R}^{4}$ is a solution to the four-dimensional autonomous system
$\frac{dU}{dx}=g(U)\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,\begin{pmatrix}u_{2}\\ -au_{1}+bu_{3}u_{1}+cu_{1}^{3}\\ u_{4}\\ -4u_{3}\end{pmatrix}$ (5)
with initial conditions $u_{3}(0)=1$ and $u_{4}(0)=0$. Then, if $u_{1},u_{2}\neq 0$ and the conditions in (3) are satisfied, the first component $u_{1}$ of $U$ is a gap soliton of the GP equation (2). An important remark is that the system (5) is conservative with conserved quantity $H=\frac{1}{2}(u_{4}^{2}+4u_{3}^{2})$. This is simply coming from the fact that in general $(u_{3}(x),u_{4}(x))=(V(x),V^{\prime}(x))$ solves the Hamiltonian system $u_{3}^{\prime}=\partial_{u_{4}}H(u_{3},u_{4})$ and $u_{4}^{\prime}=-\partial_{u_{3}}H(u_{3},u_{4})$. Now, define the set
$\gamma\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,\{(0,0,\cos(2x),-2\sin(2x)):x\in[0,\pi)\}\subset\mathbb{R}^{4},$ (6)
which represents a periodic orbit of the system (5). Looking for a gap soliton reduces to finding a solution $U:\mathbb{R}\to\mathbb{R}^{4}$ of (5) defined on all of $\mathbb{R}$, such that
@@
where $W^{u}(\gamma)$ and $W^{s}(\gamma)$ denote the unstable and stable manifolds of $\gamma$, respectively. In other words, identifying a gap soliton amounts to finding a homoclinic orbit associated with the periodic orbit $\gamma$ within a conservative four-dimensional system.
The study of connecting orbits in conservative dynamical systems is a vast and active area of research, intersecting with diverse mathematical disciplines such as algebraic topology, Morse homology, celestial mechanics, chaos theory, and the calculus of variations. Over the years, a variety of mathematical techniques have been developed to address the theoretical challenges associated with these orbits. For instance, perturbative methods within variational frameworks [11, 12, 13] and non-perturbative techniques [14, 15, 16] have been successfully employed to establish the existence of homoclinic orbits in conservative and Hamiltonian systems. Additionally, the advent of computer-assisted proofs in nonlinear analysis has greatly enriched the study of connecting orbits, employing techniques that leverage the strengths of topology, functional analysis, and scientific computing. Prominent methodologies include topological covering relations [17, 18], the Parameterization Method combined with a functional analytic setup [19, 20, 21, 22], homotopy methods [23], and interpolation-based techniques [24, 25]. These methods have been instrumental in establishing the existence of connecting orbits in the context of ordinary differential equations (ODEs) and continue to play a pivotal role in advancing the field.
In this paper, we focus on proving existence of even gap solitons in the GP equation (2) combining the aforementioned Parameterization Method for periodic orbits [26, 27, 28, 29] with computer-assisted proofs in a functional analytic sequence space framework. Let $W_{loc}^{s}(\gamma)$ denote the local stable manifold of $\gamma$. The method for obtaining its parameterization is outlined in Section 2. By restricting to the class of even solitons, we impose that $u_{1}^{\prime}(0)=0$, and hence we only need to solve for $x\geq 0$. The asymptotic condition at $x\to\infty$ for a soliton solution (3) is reinterpreted as the condition that $U(\mathsf{L})\in W_{loc}^{s}(\gamma)$, for some $\mathsf{L}\in\mathbb{R}$. From this reformulation, we deduce that the first component $u_{1}:[0,\mathsf{L}]\to\mathbb{R}$ of a solution $U:[0,\mathsf{L}]\to\mathbb{R}^{4}$ of the boundary value problem (BVP)
@@
can be extended to define an even soliton $u(x)=u_{1}(|x|)$ of the GP equation. Therefore, we will rigorously prove the existence of soliton solutions by solving the boundary value problem (7). The rest of the paper is essentially dedicated to solve rigorously the BVP (7) with the tools of computer-assisted proofs (CAPs) in nonlinear analysis (e.g. see [30, 31, 32, 33, 34] and the references therein). Specifically, once an accurate numerical approximation of a solution is available, our CAP approach guarantees a rigorous proof of the existence of a soliton that closely matches the numerical approximation, while also providing tight and precise $C^{0}$ error bounds for the discrepancy between the exact soliton and its approximation. For instance, consider Figure 1, which depicts a nontrivial approximation of an even soliton solution to the equation (2) as presented in [35]. Our method yields results of the following nature:
Theorem 1.
For the parameters $a=1.1025$, $b=0.55125$ and $c=-0.826875$, the GP equation (2) has a gap soliton solution $u:\mathbb{R}\to\mathbb{R}$.
Figure 1: The figure illustrates the different components of our method. The solution to the boundary problem is shown in blue, while the stable manifold is depicted in orange. The even extension of the soliton solution is represented in black.
It is important to emphasize that our method is general and can be applied to any set of parameters for which an accurate numerical approximation is available. To construct rigorously a parameterization of a local stable manifold of the periodic orbit $\gamma$ and subsequently to demonstrate constructively the existence of a solution of the BVP (7), we will apply the following general Newton-Kantorovich theorem.
@@
The computer-assisted part of our work relies on a numerical approximation of the soliton solution to determine $\bar{x}$. From this, we obtain a set of bounds $Y$, $Z_{1}$, and $Z_{2}$, which can be explicitly implemented in a computer program. To evaluate them numerically, we perform arithmetic operations on intervals instead of individual floating-point numbers. In this way, uncertainties propagate systematically through the computations, producing mathematically rigorous results (e.g., see [36, 31]).
The paper is structured as follows. In Section 2, we describe a computational method to obtain a parameterization of the local stable manifold $W_{loc}^{s}(\gamma)$ with rigorous error bounds. In Section 3, we introduce a constructive approach and prove the existence of solutions to the BVP (7) using Chebyshev series expansions. Finally, in Section 4, we summarize our results and provide a proof of Theorem 1, along with additional existence results.
@@ 2 Computation of a Local Stable Manifold of the Periodic Orbit
In Section 1, we reformulated the problem of proving the existence of soliton solutions as finding a solution to equation (5) that intersects the local stable manifold $W_{loc}^{s}(\gamma)$ associated with the periodic orbit $\gamma$. To explicitly characterize points on this stable manifold, we will employ the Parameterization Method [26, 27, 28], following the framework developed in [37, 29]. Since the system (5) is conservative, the periodic orbit $\gamma$ possesses two trivial Floquet exponents: one arising from the conserved quantity $H=\frac{1}{2}(u_{4}^{2}+4u_{3}^{2})$, discussed in Section 1, and another due to the shift invariance of the periodic orbit. Consequently, $\gamma$ admits at most two nontrivial Floquet exponents. For the remainder of this work, we assume that $\dim W^{u}(\gamma)=\dim W^{s}(\gamma)=1$ and denote the stable Floquet exponent of $\gamma$ by $\lambda<0$. Denote by $v:S^{1}\to\mathbb{R}^{4}$ the associated stable bundle, that is a solution of

### v2

# Computer-Assisted Proofs of Gap Solitons in Bose-Einstein Condensates

@@ Abstract
We provide a framework for turning a numerical simulation of a gap soliton in the one-dimensional Gross-Pitaevskii equation into a formal mathematical proof of its existence. These nonlinear localized solutions play a central role in understanding Bose-Einstein condensates (BECs). We reformulate the problem of proving their existence as the search for homoclinic orbits in a dynamical system. We then apply computer-assisted proof techniques to obtain verifiable conditions under which a numerically approximated trajectory corresponds to a true homoclinic orbit. This work also presents the first examples of computer-assisted proofs of gap solitons in the Gross-Pitaevskii equation on non-perturbative parameter regimes.
Key words. Bose-Einstein condensates, Gross-Pitaevskii equation, Gap solitons, Homoclinic orbits, Parameterization method for periodic orbits, Computer-Assisted Proofs
@@ 1 Introduction
A Bose-Einstein condensate (BEC) is a state of matter that forms when a collection of particles cools down to temperatures near absolute zero, causing them to lose their individual identities and merge into a single wave. The first experimental realization of BECs at ultra-cold temperatures earned the 2001 Nobel Prize in Physics, and since then, Bose-Einstein condensates have provided a platform for exploring quantum mechanics on large scales, with applications in precision measurements, quantum computing, and the modeling of complex systems such as superfluidity and optical lattices. For a comprehensive review of both experimental and theoretical developments, see [1].
Beyond their physical significance, the study of Bose-Einstein condensates offers fertile ground for advancing theoretical methods in nonlinear dynamics and partial differential equations. In this paper, we study the dynamics of a BEC using the time-dependent Gross-Pitaevskii equation, which models the BEC’s evolution in one spatial dimension:
@@
Here, $\psi(t,x)\in\mathbb{C}$ denotes the dimensionless wave function, $|\psi|^{2}$ represents the BEC density, and $V(x)$ is the external potential created by the optical lattice along the spatial domain $x\in\mathbb{R}$. The book [2] offers a comprehensive treatment of the Gross-Pitaevskii equation and nonlinear wave dynamics, combining experimental perspectives with numerical studies of Bose-Einstein condensates. It also includes an extensive bibliography covering many areas of the field.
A fundamental aspect of Bose-Einstein condensate analysis is the study of standing wave solutions to the Gross-Pitaevskii equation. These solutions take the form $\psi(t,x)=e^{-iat}u(x)$, where $u(x)$ is a real-valued function that satisfies the time-independent (GP) equation:
@@
Among the nonlinear structures admitted by the GP equation are solitons, nonlinear Bloch waves, and domain walls, each of which has been extensively studied. Our interest lies in gap solitons, a class of localized solutions. More specifically, a soliton is a real-valued function $u:\mathbb{R}\to\mathbb{R}$ that decays to zero at infinity:
@@
Chapter 19 of [2] surveys the existence of solitons under various potential types, with particular emphasis on periodic potentials. In the present work, we study the specific case $V(x)=b\cos(2x)$, which models the dynamics of a Bose-Einstein condensate in an optical lattice. To understand the setting in which gap solitons arise, we examine the linearization of the GP equation around the trivial solution, called Mathieu’s equation:
@@
Bloch theory predicts that the spectrum of this linear operator consists of bands separated by spectral gaps. The edges of these gaps are determined by solutions to (4), known as Mathieu functions. In the purely linear setting, soliton-type solutions cannot exist. When nonlinearity is introduced, however, localized modes—gap solitons—can form within the spectral gaps. These structures are characteristic of nonlinear wave systems and closely resemble gap solitons observed in nonlinear optics (e.g., see [3]).
Researchers have frequently used perturbative asymptotics and numerical simulations to investigate soliton solutions. For example, [3] applies asymptotic analytical methods to compute gap solitons in all spectral gaps of a periodic potential, showing that these solitons bifurcate from distinct band edges depending on the sign of $c$. Moreover, [4] introduces a numerical approach for calculating gap solitons in the repulsive case ($c=1$). This study shows that most solutions blow up at certain points on the real axis. Notably, the subset of non-blow-up solutions—which includes all gap solitons—forms a fractal set within the space of initial conditions. This fractal structure makes it possible to identify gap solitons over large regions in the parameter space $(a,b)\in\mathbb{R}^{2}$. While simulations and asymptotic techniques provide valuable insight, they do not yield a rigorous proof of existence. Abstract results, on the other hand, often lack any explicit description of the solution profiles.
The structure of the problem, however, makes it well-suited for a computer-assisted proof (CAP) approach. In this work, we derive verifiable conditions under which a soliton exists near a given numerical approximation. When these conditions hold, our approach guarantees a rigorous proof of existence and supplies tight, explicit $C^{0}$ error bounds for the discrepancy between the exact solution and its approximation. It is important to emphasize that our method can be applied to any set of parameters $a$, $b$ and $c$ for which an accurate numerical soliton solution is available. For example, Figure 1 shows a nontrivial numerical approximation of an even soliton solution to equation (2), as appearing in [5]. Our method provides a way to rigorously validate such approximations:
Theorem 1.
$\|u-\bar{u}\|_{\infty}\leq 8.617584260554394\cdot 10^{-6},$
where $\bar{u}$ is a numerical approximation of the solution illustrated in Figure 1.
Figure 1: The figure shows a validated soliton solution of the Gross-Pitaevskii (GP) equation with parameters $a=1.1025$, $b=0.55125$, and $c=-0.826875$. It also depicts the main elements of our approach: the solution to the boundary-value problem (blue), the stable manifold (orange), and the even extension of the soliton (black).
[…]
with initial conditions $u_{3}(0)=1$ and $u_{4}(0)=0$. Then, if $u_{1},u_{2}\neq 0$ and the conditions in (3) are satisfied, the first component $u_{1}$ of $U$ is a gap soliton of the GP equation (2). We emphasize that the system (5) is not conservative, though it possesses the conserved quantity $H=\frac{1}{2}(u_{4}^{2}+4u_{3}^{2})$. This comes from the fact that $(u_{3}(x),u_{4}(x))=(V(x),V^{\prime}(x))$ solves the Hamiltonian system $u_{3}^{\prime}=\partial_{u_{4}}H(u_{3},u_{4})$, $u_{4}^{\prime}=-\partial_{u_{3}}H(u_{3},u_{4})$. Now, define the set
[…]
which represents a periodic orbit of the system (5). Looking for a gap soliton reduces to finding a solution $U:\mathbb{R}\rightarrow\mathbb{R}^{4}$ of (5) defined on all of $\mathbb{R}$, such that
@@
where $W^{u}(\gamma)$ and $W^{s}(\gamma)$ denote the unstable and stable manifolds of $\gamma$, respectively. In other words, identifying a gap soliton amounts to finding a homoclinic orbit associated with the periodic orbit $\gamma$ in the four-dimensional system (5) .
The study of connecting orbits in dynamical systems is a vast and active area of research, intersecting with diverse mathematical disciplines such as algebraic topology, Morse homology, celestial mechanics, chaos theory, and the calculus of variations. Over the years, a variety of mathematical techniques have been developed to address the theoretical challenges associated with these orbits. For instance, perturbative methods within variational frameworks [6, 7, 8] and non-perturbative techniques [9, 10, 11] have been successfully employed to establish the existence of homoclinic orbits in conservative and Hamiltonian systems. Additionally, the advent of computer-assisted proofs in nonlinear analysis has greatly enriched the study of connecting orbits, employing techniques that leverage the strengths of topology, functional analysis, and scientific computing. Prominent methodologies include topological covering relations [12, 13], the Parameterization Method combined with a functional analytic setup [14, 15, 16, 17], homotopy methods [18], and interpolation-based techniques [19, 20]. These methods have been instrumental in establishing the existence of connecting orbits in the context of ordinary differential equations (ODEs) and continue to play a pivotal role in advancing the field.
Let $W_{loc}^{s}(\gamma)$ denote the local stable manifold of $\gamma$. By restricting to the class of even solitons, we impose that $u_{1}^{\prime}(0)=0$, and hence, we only need to solve for $x\geq 0$. The asymptotic condition at $x\to\infty$ for a soliton solution (3) is reinterpreted as the condition that $U(\mathsf{L})\in W_{loc}^{s}(\gamma)$, for some $\mathsf{L}\in\mathbb{R}$. From this reformulation, we deduce that the first component $u_{1}:[0,\mathsf{L}]\to\mathbb{R}$ of a solution $U:[0,\mathsf{L}]\to\mathbb{R}^{4}$ of the boundary-value problem (BVP)
@@
can be extended to define an even soliton $u(x)=u_{1}(|x|)$ of the GP equation. To obtain an explicit boundary condition on the stable manifold, we use the Parameterization Method for periodic orbits [21, 22, 23, 24].
The rest of the paper is dedicated to solving the boundary-value problem (7). We reformulate both the construction of a local stable manifold of the periodic orbit $\gamma$ and the boundary-value problem as zero-finding problems in infinite-dimensional Banach spaces. In both cases, we apply tools from computer-assisted proofs in nonlinear analysis [25, 26, 27, 28, 29] to prove the existence of solutions near approximate zeros of the corresponding maps. Our approach is based on the Newton–Kantorovich-type theorem stated below.
@@
In our setting, $\bar{x}$ denotes a numerical approximate solution. The quantities $Y$, $Z_{1}$, and $Z_{2}$ correspond to bounds on the norms of elements in certain infinite-dimensional Banach spaces. For any given set of parameters, we derive these bounds in a form that can be rigorously evaluated by a computer. This derivation constitutes the main technical part in our paper.
While other approaches to studying the existence of gap solitons exist, our main contribution is to provide explicit conditions that guarantee the existence of a true solution near a numerical approximation. Moreover, the examples we present constitute the first computer-assisted proofs of soliton existence in the Gross–Pitaevskii equation in non-perturbative parameter regimes.
Our paper is structured as follows. In Section 2, we describe a computational method to obtain a parameterization of the local stable manifold $W_{loc}^{s}(\gamma)$ with rigorous error bounds. In Section 3, we introduce a constructive approach and prove the existence of solutions to the BVP (7) using Chebyshev series expansions. Finally, in Section 4, we provide examples of our method, including a computer-assisted proof of Theorem 1.
@@ 2 Computation of a Local Stable Manifold of the Periodic Orbit
In Section 1, we reformulated the problem of proving the existence of soliton solutions as finding a solution to equation (5) that intersects the local stable manifold $W_{loc}^{s}(\gamma)$ associated with the periodic orbit $\gamma$. To explicitly characterize points on this stable manifold, we will employ the Parameterization Method [21, 22, 23], following the framework developed in [30, 24]. The periodic orbit $\gamma$ possesses two trivial Floquet exponents: one arising from the conserved quantity $H=\frac{1}{2}(u_{4}^{2}+4u_{3}^{2})$, discussed in Section 1, and another due to the shift invariance of the periodic orbit. Consequently, $\gamma$ admits at most two nontrivial Floquet exponents. For the remainder of this work, we assume that $\dim W^{u}(\gamma)=\dim W^{s}(\gamma)=1$ and denote the stable Floquet exponent of $\gamma$ by $\lambda<0$. Denote by $v:S^{1}\to\mathbb{R}^{4}$ the associated stable bundle, that is a solution of